# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ IV-A2 Agent-Agent Authentication Threats
As LM agents increasingly collaborate to tackle complex tasks through horizontal, vertical, or hybrid coordination, authentication between LM agents becomes critical to LM agent networks. A primary threat in such scenarios is agent spoofing attacks, where adversaries compromise or impersonate legitimate agents within the collaboration workflow. For instance, in horizontal or vertical collaboration scenarios, attackers may compromise LM agents, converting them into malicious entities, or impersonate legitimate agents to disrupt collaborative LM inference processes, thereby undermining the accuracy and reliability of the collaboration system and jeopardizing the task’s outcome, as illustrated in Fig. 21(b). Additionally, collusive LM agents can sabotage the consensus-building process required for task execution, introducing misinformation or bias that disrupts the overall functionality and trustworthiness of the collaborative system.
@@ IV-A3 Agent-TPSP Authentication Threats
LM agents often rely on external tools, web APIs, and third-party service providers (TPSPs) to execute diverse tasks, making secure and reliable authentication between agents and TPSPs critical. This process involves mutual verification: (i) TPSPs should confirm the agent’s legitimacy and authorization, typically via a valid API key or token; (ii) agents should authenticate TPSPs to ensure they interact with trusted and legitimate service endpoints. Authentication threats in this context primarily target the communication channel between LM agents and TPSPs, exposing them to interception-based attacks such as man-in-the-middle (MITM) and session hijacking. For instance, an adversary could intercept agent-TPSP communications, altering responses or injecting erroneous data, leading to unexpected outputs for users, as illustrated in Fig. 21(c). Additionally, improper API key or token management can exacerbate risks and result in unauthorized service misuse.
@@ IV-A4 Countermeasures to Authentication Threats in LM Agent Networks
While existing technologies such as end-to-end encryption, certificates, MFA, and key management are effective in mitigating authentication threats in LM agent services, their effectiveness diminishes in the face of the growing complexity of LM agent collaboration and the evolving sophistication of attacks driven by LMs. Moreover, the rapid development of LM agents has created a dual-edged dynamic, where both attacks and defenses are increasingly automated, escalating the risks and complexities of maintaining security. As a result, it is crucial to develop adaptive and collaborative resilient authentication mechanisms to ensure the security and reliability of LM agent systems in complex, dynamic environments.
@@ IV-B3 Countermeasures to DoS Attacks in LM Agent Networks
  • Malicious inputs detection and filtering technologies can effectively identify abnormal inputs, such as sponge examples, by comparing their energy consumption with that of natural inputs [121]. Specifically, prior to deploying LM agents, an analysis of natural inputs can be conducted to measure the time or energy consumption during inference. The defender can then establish a cutoff threshold to limit the maximum energy consumption for each inference, thus mitigating the impact of malicious inputs (e.g., sponge examples).
@@
IV-C Hallucination in LM Agent Networks
Figure 23: Taxonomy of hallucinations in the context of LM agents.
Decision-making reliability is crucial for practically deploying LM agent services. The hallucination phenomenon in LM agents refers to the generation of erroneous or illogical outputs that deviate from user inputs, generated context, or real-world conditions, posing a significant risk to service reliability. In multi-agent scenarios, hallucinations can be exacerbated through agent interactions, amplifying and propagating errors across the system. According to [34], we categorize the hallucination within the context of LM agents from the following four perspectives, as illustrated in Fig. 23.
IV-C1 Hallucination from Data and Knowledge
The primary cause of hallucinations in LM agents stems from the biased or imperfect nature of the training data and knowledge employed for content generation. Specifically, i) annotation irrelevance in the collected dataset can lead to hallucinations in LM agents. For LLMs, source-reference divergence occurs due to heuristic data collection, implying that the selected sources may lack sufficient information for content generation [122]. For LVMs, a substantial amount of instruction data is synthesized using LLMs. However, generated instructions may not accurately correspond to the content depicted in the images due to the unreliability of generative models [123]. ii) The presence of distribution imbalance in the training data is another cause for hallucinations. The training dataset contains significantly more positive samples than negative ones, biasing LMs towards responding “yes” in binary discrimination scenarios [123, 124]. iii) Duplicates within the training data may not be adequately filtered during pre-processing. Consequently, the trained LMs might tend to produce content that mimics the patterns in these duplicates [125]. iv) The training data and the knowledge, which can be inherently hallucinatory due to bias, outdated content, or fabrication [126]. LMs agents trained on these data are susceptible to replicate or even amplify hallucinations. Anomalous retrieval from the knowledge base can also introduce biases into the dependent knowledge during content generation process, thereby leading to the emergence of hallucinations in the generated content.
IV-C2 Hallucination from Training and Inference
Parikh et al. [127] demonstrate that hallucinations can occur even when the training dataset is nearly unbiased, and another key factor contributing to hallucinations is the presence of imperfections during LM training and inference processes. Specifically, i) defective encoder and decoder components in LM agents can result in hallucinations. The flawed representation learning in LM training increases the unexpected nature of the generation, thereby increasing the likelihood of hallucinated content [127]. Moreover, the sampling-based top-p decoding strategy has been demonstrated to be positively correlated with increased hallucination of generated content due to its inherent randomness [128]. ii) Parametric knowledge refers to LMs’ ability to utilize model parameters to memorize relevant information (e.g., knowledge) from training data, which can assist in improving the performance of downstream tasks [112]. Recent research has found that LMs prioritize parametric knowledge over user input during content generation [129], implying that biased parametric knowledge can result in a large amount of hallucinated information. iii) A problematic alignment process can also lead to hallucinations in LM agents. If the necessary prior knowledge is not adequately acquired during LM pre-training phase, the agent may produce hallucinated responses. Additionally, sycophancy, where the agent generates answers that align more with the user’s viewpoint than with accurate and credible information, can also contribute to hallucinations [130]. iv) The token generation strategy of LM agents may give rise to a phenomenon known as hallucination snowballing, wherein LM agent tends to persist in early errors for self-consistency rather than correcting them [131].
IV-C3 Hallucination from Multi-agent Interaction
In multi-agent environments, interactions between LM agents can give rise to new hallucination threats. Specifically, i) conflicts between multiple LM agents may result in hallucinations in final responses. Due to the diverse objectives, strategies, and knowledge of individual agents, conflicting viewpoints or misinformation can inadvertently influence the final output when these agents engage in communication or collaboration. The potential adversarial manipulation may exacerbate the occurrence of conflicts, thereby leading to more severe hallucinations. ii) In multi-agent systems, the occurrence of a single agent’s hallucination can initiate a cascading effect, meaning that misinformation from one agent may be accepted and further propagated by others within the LM agent network [44, 68], especially under vertical collaboration paradigms. Addressing hallucinations in LM agent networks requires not only correcting inaccuracies at the individual agent level but also effectively managing the exchange of information among agents to prevent the spread of hallucinations.
IV-C4 Countermeasures to Hallucination in LM Agent Networks
Researchers have developed a suite of countermeasures targeting hallucination threats within LM agents.
  • Data sanitization represents an intuitive approach to mitigate hallucinations by excluding as much unreliable data as possible from training dataset. The manually-revised controlled dataset named ToTTo is validated to improve the factuality of results on table-to-text generation tasks [127]. Additionally, RefinedWeb is an automatically filtered and deduplicated web dataset, which can mitigate hallucinations and improve LLM performance [126].
  • Reinforcement learning prevents LMs from generating hallucinatory information by introducing external feedbacks and reward mechanisms to constrain LM agents’ behavior, facilitating LM agents to refrain from QA beyond their capability instead of fabricating untruthful responses [132, 77]. For instance, GPT-4 collects the unfactual data annotated by users and generates synthetic closed-domain hallucinated synthetic data to train the reward model, reducing the model’s tendency to hallucinate [7].
  • Hallucination detection is a promising strategy which detects hallucinations in generated content and subsequently regenerates samples to eliminate inaccuracies. Techniques such as SelfcheckGPT [133] and INSIDE [134] involve assessing the consistency across multiple generated responses to identify the occurrence of hallucinations.
  • Truthful external knowledge & RAG techniques can facilitate truthful content generation [58, 77]. They concatenate external knowledge with user input to establish contextual knowledge, which is then used to query LM agents to reduce hallucinations in generated responses [135].
  • Instruction tuning is an effective approach to mitigate hallucinations, by optimizing instructions or fine-tuning LMs on a robust instruction dataset. SYNTRA [136] evaluates hallucinations on synthesis tasks and optimizes transferred instructions to reduce hallucinations on downstream tasks. Moreover, Liu et al. [123] mitigate hallucinations in LVMs by constructing a robust instruction dataset named LRV-Instuction and fine-tuning LVMs on it.
  • Post-processing techniques can help correct hallucinations in LM responses [137, 138]. For instance, RARR [137] utilizes search engines to gather relevant external knowledge and rectifies generated content from multiple perspectives. Meanwhile, LURE [138] reduces hallucinations by correcting object-level inaccuracies during the post-processing phase.
  • Model architecture optimization is effective to mitigate hallucinations. Multi-branch decoder [139] and uncertainty-aware decoder [140] are two examples of modified decoder structures to mitigate hallucinations.
  • Meta programming framework can effectively guide and promote the collaboration among LM agents, thereby reducing hallucinations in complex tasks [68].
IV-D Adversarial Attacks in LM Agent Networks
@@
For traditional AI models, an adversarial attack involves an adversary subtly manipulating the input by adding imperceptible perturbations, causing the AI model to produce outputs that deviate from the expected results. In the realm of LM agents, there are two types of adversarial attacks: adversarial input attacks and prompt hacking attacks, as depicted in Fig. 24.
IV-D1 Adversarial Input Attack
@@
Adversarial input attacks are analogous to traditional generative adversarial attacks, where the attacker degrades the accuracy of generated content in LM agents by manipulating input instructions, as depicted in Fig. 24(a). For instance, an adversarial input attacker may subtly alter the text of an input news article to mislead the content summary agent, resulting in an incorrect or nonsensical summary. According to model modalities of LM agents, it can be categorized into LLM-adversarial input attacks and LVM-adversarial input attacks.
@@
  • LLM-adversarial input attacks refers to the adversarial attack on LLMs that perturbs the input text into the semantic equivalent adversarial text, aiming to confuse LLMs to generate erroneous contents or adversary-desired contents. Zhuo et al. [32] demonstrate that CODEX, a pretrained LLM capable of executing prompt-based semantic parsing on codes, is vulnerable to adversarial inputs, especially those generated by sentence-level perturbations. Xu et al. [141] propose PromptAttack, which converts adversarial textual attacks into an attack prompt, thereby causing the target LLM to generate adversarial examples to fool itself. Shi et al. [142] show that LLMs can be easily distracted by adding a small amount of irrelevant information to inputs, which significantly reduces the accuracy of their responses. Liu et al. [143] highlight that current LLMs do not involve social interactions during their training process. Consequently, the oversight leads to poor generalization ability in unfamiliar cases and a lack of robustness against adversarial attacks.
@@
  • LVM-adversarial input attack refers to the adversarial attack on LVMs (e.g., CLIP and MiniGPT-4) that adds carefully crafted imperceptible adversarial perturbations to vulnerable modalities in the joint input prompts (i.e., visual, textual, or both), thereby compromising corresponding outputs. Du et al. [144] introduce a new adversarial attack named auto-attack on text-to-image models by adding small perturbations generated by gradient optimization to text prompts, resulting in the fusion of main subjects with unrelated categories or even their complete disappearance in generated images. The transferability of adversarial input attacks in LVMs has been extensively studied [145, 146]. Wang et al. [145] introduce the transferable multi-modal (TMM) attack, which integrates attention-directed feature perturbation and orthogonal-guided feature heterogenization to generate transferable adversarial examples against LVMs. Luo et al. [146] develop the cross-prompt attack (CroPA), which leverages the learnable prompts to generate transferable adversarial images across LVMs. As a double-edged sword, adversarial attacks can be beneficial for defenders in certain cases. For instance, Liang et al. [147] leverage these attacks in the context of diffusion models to inhibit the learning, imitation, and replication of images by such models, thereby protecting intellectual property rights of artists.
IV-D2 Prompt Hacking Attacks
@@
As depicted in Fig. 24(b), prompt hacking attacks involve using specifically crafted input instructions to bypass the security constraints of LM agents, thereby generating harmful contents. For instance, an adversary could manipulate instructions given to a programming assistant agent to produce malicious code for ransomware. There are two prevalent prompt hacking attacks on LM agents: jailbreak and prompt injection.
  • Jailbreak: LM agents typically achieve inherent predefined rule restrictions through model alignment techniques, thereby preventing the generation of harmful or malicious content. Jailbreak refers to adversaries designing particular prompts that exploit model vulnerabilities, bypassing the content generation rules, and enabling LM agents to generate harmful or malicious content [148, 149, 150, 151]. Yu et al. [148] evaluate hundreds of jailbreak prompts on GPT-3.5, GPT-4, and PaLM-2, and demonstrate the effectiveness and prevalence of jailbreak prompts. Yang et al. [149] propose an automated jailbreak attack framework named SneakyPrompt, which successfully jailbreaks DALL-E 2. SneakyPrompt effectively bypasses the safety filters and enable generation of not-safe-for-work (NSFW) images. Shen et al. [150] perform a comprehensive measurement study on in-the-wild jailbreak prompts, identifying two long-term jailbreak prompts can achieve 99% attack success rates on GPT-3.5 and GPT-4. Deng et al. [151] introduce an end-to-end jailbreak framework named MASTERKEY, which utilizes the reverse engineering to uncover the internal jailbreak defense mechanisms of LLMs and leverages a fine-tuned LLM to automatically generate jailbreak prompts. In practical environments, jailbreak poses a more significant threat to embodied LM agents. Zhang et al. [152] identify three risks associated with embodied AI agents based on LLMs: cascading vulnerability propagation, cross-domain safety misalignment, and conceptual deception challenges. The first risk involves exploiting inherent jailbreak vulnerabilities in LLMs to induce harmful behaviors. The second arises from the mismatch between the language output and action output spaces of embodied LLM agents, allowing for the rejection of harmful behavior linguistically while still executing harmful actions. The underlying reason of this risk may stem from the inherent differences in aligned data. Currently, there is a wealth of aligned data within the language space, including a significant amount of ethically sound texts. In contrast, aligned data oriented towards action outputs is relatively scarce. The third risk stems from the current lack of reasoning capabilities in embodied LLM agents, making them susceptible to conceptual deception by attackers, which can result in seemingly harmless actions leading to malicious outcomes. For instance, an embodied LLM agent may decline a direct command to “poison the person”, but follow a series of inducible instructions that lead to the same result, such as “put the poison in the person’s mouth”.
  • Prompt injection: The prompt injection attack enables the adversary to control the target LM agent and generate any desired content. The attack is performed by manipulating user inputs to create meticulously crafted prompts, so that LM agents are unable to distinguish between developer’s original instructions and user inputs. These prompts then hijack the agent’s intended task, causing the agent’s outputs to deviate from expected behaviors [153, 154, 36]. Toyer et al. [153] collect a dataset comprising prompt injection attacks and prompt-based injection defenses from players of an online game called Tensor Trust. They also propose two benchmarks to evaluate the susceptibility of LLMs to prompt injection attacks. Liu et al. propose HOUYI [36], an innovative black-box prompt injection attack inspired by traditional web injection attacks. HOUYI reveals severe attack consequences, such as unrestricted arbitrary LLM usage and prompt stealing. Greshake et al. introduce the concept of indirect prompt injection attack [154], which leverages LM agents to inject crafted prompts into the data retrieved at inference time. Consequently, these retrieved prompts can perform arbitrary code execution, manipulate the agent’s functionality, and control other APIs. Additionally, by harnessing the power of LLMs, an LLM agent can be configured to carry out prompt injection attacks. Ning et al. propose CheatAgent [155], a novel LLM-based attack framework that generates adversarial perturbations on input prompts to mislead black-box LLM-powered recommender systems.
IV-D3 Countermeasures to Adversarial Attacks in LM Agent Networks
@@
Existing countermeasures to adversarial attacks to secure LM agents involve adversarial training, input/output filtering, robust optimization, and auditing & red teaming.
@@
  • Adversarial training aims to enhance LM’s robustness in the input space by incorporating adversarial examples into the training data. Bespalov et al. [156] demonstrate that basic adversarial training can significantly improve the robustness of toxicity language predictors. Cheng et al. introduce AdvAug [157], a novel adversarial augmentation method for neural machine translation (NMT) to enhance translation performance.
@@
  • Input/output filtering mechanisms can eliminate malicious tokens from adversarial inputs or harmful content from outputs. Kumar et al. [158] propose the erase-and-check method, which leverages another LLM as a safety filter to remove malicious tokens in the user input. Phute et al. [159] present an LLM self-examination defense approach, where an extra LLM is utilized to evaluate whether the responses are generated by adversarial prompts. Zeng et al. [160] propose AutoDefense, a multi-agent framework to defend against jailbreak attacks by filtering harmful LLM responses without impacting user inputs. AutoDefense divides the defense task into sub-tasks, leveraging LLM agents based on AutoGen [13] to handle each part independently. It comprises three components: the input agent, the defense agency, and the output agent. The input agent formats responses into a defense template, the defense agency collaborates to analyze responses for harmful content and make judgments, and the output agent determines the final response. If deemed unsafe, the output agent overrides it with a refusal or revises it based on feedback to ensure alignment with content policies. Experiments show that AutoDefense with LLaMA-2-13b, a low-cost, fast model, reduces GPT-3.5’s attack success rate (ASR) from 55.74% to 7.95%, achieving 92.91% defense accuracy.
@@
  • Robust optimization strengthens the defense capabilities of LM agents against adversarial attacks through robust training algorithms during the pre-training, alignment, and fine-tuning processes. Shen et al. [161] propose a dynamic attention method that mitigates the impact of adversarial attacks by masking or reducing the attention values assigned to adversarial tokens.
@@
  • Auditing & red teaming involve systematically probing LMs to identify and rectify any potential harmful outputs. Jones et al. [162] introduce ARCA, a discrete optimization algorithm designed to audit LLMs, which can automatically detect derogatory completions about celebrities, thus providing a valuable tool for uncovering model vulnerabilities before deployment. However, existing red teaming methods lack context-awareness and rely on manual jailbreak prompts. To address this, Xu et al. [163] propose RedAgent, a multi-agent LLM system that generates context-aware jailbreak prompts using a coherent set of jailbreak strategies. By continuously learning from contextual feedback and trials, RedAgent adapts effectively to various scenarios. Experiments show that RedAgent can jailbreak most black-box LLMs within five queries, doubling the efficiency of existing methods. Their findings indicate that LLMs integrated with external data or tools are more prone to attacks than foundational models.
IV-E Poisoning & Backdoor Attacks in LM Agent Networks

### v3

@@ IV-A2 Agent-Agent Authentication Threats
As LM agents increasingly collaborate to tackle complex tasks through horizontal, vertical, or hybrid coordination, authentication between agents becomes critical. One major threat is agent spoofing, where adversaries compromise or impersonate legitimate agents within the collaboration workflow. In both horizontal and vertical collaboration scenarios, attackers may compromise LM agents, converting them into malicious entities or impersonating legitimate ones to disrupt collaborative inference processes. As illustrated in Fig. 25(b), it undermines accuracy and reliability of the collaboration system and jeopardizes task’s outcome. Additionally, collusive LM agents can sabotage consensus-building, injecting misinformation or bias that disrupts the overall functionality and trustworthiness of the collaborative system.
@@ IV-A3 Agent-TPSP Authentication Threats
LM agents often rely on external tools, web APIs, and third-party service providers (TPSPs) to execute diverse tasks, making secure and reliable authentication between agents and TPSPs critical. This process involves mutual verification: (i) TPSPs should confirm the agents’ legitimacy and authorization, typically via a valid API key or token; (ii) agents should authenticate TPSPs to ensure they interact with trusted and legitimate service endpoints. Authentication threats in this context primarily target the communication channel between LM agents and TPSPs, exposing them to interception-based attacks such as man-in-the-middle (MITM) and session hijacking. For instance, an adversary could intercept agent-TPSP communications, altering responses or injecting erroneous data, leading to unexpected outputs for users, as illustrated in Fig. 25(c). Additionally, improper API key or token management can exacerbate risks and result in unauthorized service misuse.
@@ IV-A4 Countermeasures to Authentication Threats in LM Agent Networks
While existing technologies such as end-to-end encryption, certificates, MFA, and key management are effective in mitigating authentication threats in LM agent services, their effectiveness diminishes in the face of the growing complexity of LM agent collaboration and the evolving sophistication of attacks driven by LMs. Moreover, the rapid development of LM agents has created a dual-edged dynamic, where both attacks and defenses are increasingly automated, escalating the risks and complexities of maintaining security. As a result, it is crucial to develop adaptive and collaborative resilient authentication mechanisms to guarantee the security and reliability of LM agent systems in complex, dynamic environments.
@@ IV-B3 Countermeasures to DoS Attacks in LM Agent Networks
  • Malicious input detection and filtering technologies can effectively identify abnormal inputs, such as sponge examples, by comparing their energy consumption with that of natural inputs [125]. Specifically, prior to deploying LM agents, an analysis of natural inputs can be conducted to assess the time and energy consumption during inference. Based on this analysis, the defender can set a cutoff threshold to limit the maximum energy consumption for each inference, thus mitigating the impact of malicious inputs (e.g., sponge examples).
@@
IV-C Adversarial Attacks in LM Agent Networks
@@
For traditional AI models, an adversarial attack involves an adversary subtly manipulating the input by injecting imperceptible perturbations, causing the AI model to generate outputs that deviate from the expected results. In the realm of LM agents, there exist two types of adversarial attacks: adversarial input attacks and prompt hacking attacks, as depicted in Fig. 27.
IV-C1 Adversarial Input Attack
@@
Adversarial input attacks are analogous to traditional generative adversarial attacks, where the attacker degrades the accuracy of generated content in LM agents by manipulating input instructions, as depicted in Fig. 27(a). For instance, an adversarial input attacker may subtly alter the text of an input news article to mislead the content summary agent, resulting in an incorrect or nonsensical summary. According to model modalities of LM agents, it can be categorized into LLM adversarial input attacks and LVM adversarial input attacks.
@@
a) LLM adversarial input attacks involve perturbing input text into semantically equivalent adversarial forms to mislead LLMs into generating erroneous or adversary-desired outputs. Zhuo et al. [26] demonstrate that CODEX, a pretrained LLM for prompt-based semantic code parsing, is vulnerable to adversarial inputs, especially those generated through sentence-level perturbations. Xu et al. [126] introduce PromptAttack, which converts adversarial textual attacks into attack prompts, causing LLMs to generate adversarial examples that deceive themselves. Similarly, Shi et al. [127] demonstrate that injecting small amounts of irrelevant information can significantly degrade an LLM’s accuracy. Liu et al. [128] further highlight a key limitation: current LLMs lack exposure to social interactions during training, leading to poor generalization in unfamiliar scenarios and reduced robustness against adversarial attacks.
b) LVM adversarial input attacks target LVMs (e.g., CLIP and MiniGPT-4) by adding carefully crafted imperceptible adversarial perturbations to vulnerable modalities in joint input prompts (i.e., visual, textual, or both), thereby compromising output integrity. Du et al. [129] introduce an adversarial attack named auto-attack on text-to-image models by adding small gradient-based perturbations to text prompts, causing the fusion of main subjects with unrelated categories or even their complete disappearance in generated images.
@@
The transferability of adversarial input attacks in LVMs has been extensively studied [130, 131]. Wang et al. [130] introduce the transferable multi-modal (TMM) attack, which combines attention-directed feature perturbation and orthogonal-guided feature heterogenization to generate transferable adversarial examples against LVMs. Similarly, Luo et al. [131] present the cross-prompt attack (CroPA), which leverages learnable prompts to craft adversarial images transferable across different LVMs. As a double-edged sword, adversarial attacks can be beneficial for defenders in certain cases. For instance, Liang et al. [132] leverage adversarial techniques in diffusion models to prevent unauthorized learning, imitation, and replication of images, thereby safeguarding artists’ intellectual property rights.
IV-C2 Prompt Hacking Attacks
@@
As depicted in Fig. 27(b), prompt hacking involves using specifically crafted input instructions to bypass security constraints of LM agents, thereby generating harmful contents. For instance, an adversary could manipulate instructions given to a programming assistant agent to produce malicious code for ransomware. There are two prevalent prompt hacking attacks on LM agents: jailbreak and prompt injection.
a) Jailbreak: LM agents typically enforce predefined rule restrictions through model alignment techniques, preventing the generation of harmful or malicious content. However, jailbreaking occurs when adversaries craft particular prompts that exploit model vulnerabilities, bypassing content generation rules and enabling the generation of harmful outputs [133, 134, 135, 136]. Yu et al. [133] evaluate hundreds of jailbreak prompts on GPT-3.5, GPT-4, and PaLM-2, demonstrating their effectiveness and widespread impact. Yang et al. [134] propose an automated jailbreak framework named SneakyPrompt, which successfully bypasses DALL-E 2’s safety filters to generate not-safe-for-work (NSFW) images. Shen et al. [135] perform a comprehensive measurement study on in-the-wild jailbreak prompts, identifying two long-term jailbreak prompts that achieve a 99% attack success rate (ASR) on GPT-3.5 and GPT-4. Deng et al. [136] introduce an end-to-end jailbreak framework named MASTERKEY, which reverse-engineers LLM defense mechanisms and leverages fine-tuned models to generate jailbreak prompts automatically.
Jailbreaking poses an even greater risk to embodied LM agents in real-world applications. Zhang et al. [137] identify three risks associated with embodied AI agents based on LLMs.
  • Cascading vulnerability propagation: Attackers exploit jailbreak vulnerabilities in LLMs to induce harmful behaviors.
  • Cross-domain safety misalignment: A mismatch between language and action output spaces enables an agent to reject harmful behavior linguistically while still executing it in practice. This likely stems from the abundance of aligned textual data within the language space, compared to the scarcity of aligned action-oriented data.
  • Conceptual deception: Limited reasoning capabilities make embodied LLM agents vulnerable to indirect manipulation, which can result in seemingly harmless actions but actually malicious outcomes. For instance, an embodied LLM agent may decline a direct command to “poison the person”, but still follow step-by-step instructions leading to the same outcome, such as “put the poison in the person’s mouth”.
b) Prompt injection: Prompt injection attacks enable adversaries to manipulate target LM agents into generating unintended content by crafting deceptive user inputs. These attacks exploit the agent’s inability to distinguish between a developer’s original instructions and manipulated prompts, effectively hijacking its intended functionality [138, 139, 30], leading to deviations in the agent’s output from expected behaviors. Toyer et al. [138] construct a dataset comprising prompt injection attacks and corresponding defenses from an online game named Tensor Trust. Additionally, they introduce two benchmarks for assessing LLM vulnerability to prompt injection threats. Liu et al. introduce HOUYI [30], a black-box prompt injection attack inspired by traditional web injection techniques, which enables extensive misuse of LLMs and prompt theft. Greshake et al. introduce the indirect prompt injection attack [139], where LM agents inject crafted prompts into data retrieved at inference time. These manipulated prompts can execute arbitrary code, alter agent behavior, and take control of external APIs.
Beyond external attacks, LM agents themselves can be configured to execute prompt injection. Ning et al. propose CheatAgent [140], an LLM-based attack framework that generates adversarial perturbations in input prompts to mislead black-box LLM-powered recommender systems.
IV-C3 Countermeasures to Adversarial Attacks in LM Agent Networks
@@
Existing countermeasures against adversarial attacks on LM agents include adversarial training, input/output filtering, robust optimization, and auditing & red teaming.
@@
  • Adversarial training aims to enhance an LM’s robustness in the input space by integrating adversarial examples into the training process. Bespalov et al. [141] demonstrate that basic adversarial training substantially improves the resilience of toxicity language predictors. Cheng et al. introduce AdvAug [142], an adversarial augmentation technique designed to boost neural machine translation (NMT) performance.
@@
  • Input/output filtering mechanisms can remove malicious tokens from adversarial inputs or harmful content from outputs. Kumar et al. [143] introduce the erase-and-check method, which employs an auxiliary LLM as a safety filter to eliminate malicious tokens from user inputs. Phute et al. [144] present an LLM self-examination defense approach, where an additional LLM assesses whether responses are generated from adversarial prompts. Zeng et al. [145] propose AutoDefense, a multi-agent framework that mitigates jailbreak attacks by filtering harmful responses without altering user inputs. AutoDefense divides the defense task into sub-tasks, utilizing LLM agents based on AutoGen [9] to handle each part independently, which consists of three components: an input agent, a defense agency, and an output agent. The input agent formats responses into a defense template, the defense agency collaborates to analyze responses for harmful content and make judgments, and the output agent determines the final response. If deemed unsafe, the output agent overrides it with a refusal or revises it based on feedback to ensure compliance with content policies. Experiments show that AutoDefense, implemented with LLaMA-2-13b, reduces GPT-3.5’s ASR from 55.74% to 7.95%, achieving 92.91% defense accuracy.
@@
  • Robust optimization reinforces the defense capabilities of LM agents against adversarial attacks through robust training algorithms during pre-training, alignment, and fine-tuning processes. Shen et al. [146] propose a dynamic attention method that mitigates adversarial threats by masking or reducing attention values assigned to adversarial tokens.
@@
  • Auditing & red teaming involve systematically probing LMs to identify and rectify potential harmful outputs. Jones et al. [147] introduce ARCA, a discrete optimization algorithm for auditing LLMs, capable of automatically detecting derogatory completions about celebrities, providing a valuable tool for uncovering model vulnerabilities before deployment. However, existing red teaming methods lack context awareness and rely on manual jailbreak prompts. To address this, Xu et al. [148] propose RedAgent, a multi-agent LLM system that generates context-aware jailbreak prompts using a coherent set of jailbreak strategies. By continuously learning from feedback and trials, RedAgent adapts dynamically to various scenarios. Experimental results show that RedAgent successfully jailbreaks most black-box LLMs in fewer than five queries, achieving twice the efficiency of current approaches. Additionally, their findings suggest that LLMs integrated with external data or tools are more vulnerable to attacks than standalone foundational models.
IV-D Poisoning & Backdoor Attacks in LM Agent Networks